1. xy = 3y + 3 3. (x + 1)y - (2x + 3) y = 0
Answer #1

Se utim (a) xy xy = 34+3 se du 34 +3 → dy 12 3 32 ły 3 x which is of type de la + py of x. Here, py = , where P and Q are f⑧ (x+1) d. -(43)g - 0 가 dy = (21439 da 생 (2143) da (4) dy Cortegrating ( 23) da () my) = 21 +2 + 1) (+1) da j (2+2) + 1 (+1)(2717) ܩܘ (22+) Answer.

If any doubt please comment